# Stress out of charmonia

### v1

@@ Abstract
We investigate the gravitational form factors of charmonium. Our method is based on a Hamiltonian formalism on the light front known as basis light-front quantization. The charmonium mass spectrum and light-front wave functions were obtained from diagonalizing an effective Hamiltonian that incorporates confinement from holographic QCD and one-gluon exchange interaction from light-front QCD. We proposed a quantum many-body approach to construct the hadronic matrix elements of the energy momentum tensor $T^{++}$ and $T^{+-}$, which are used to extract the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$. The obtained form factors satisfy the known constraints, e.g. von Laue condition. From these quantities, we also extract the energy, pressure and invariant mass squared distributions of the system. We find that hadrons are multi-layer systems. We also find that there is a tachyonic core within charmonium.
@@ I Introduction
The year 2023 marked the 50 year anniversary of quantum chromodynamics (QCD). However, our understanding of the strong interaction remains incomplete [1]. One of the challenges is the strong-coupling nature of QCD. As such, the strong force within hadrons is dramatically different from what the QCD Lagrangian naïvely tells us, in contrast to weakly coupled systems, e.g. the hydrogen atom. The primary operators that describe the distribution of stress within a system are the energy-momentum tensor (EMT) $T^{\mu\nu}$ [2]. These operators also dictate how the system gravitates, and conversely, how it responds to the gravitational field [3, 175]. The access to the hadronic matrix elements (HME) of the EMT has been a hot topic in recent literature. Kumano et al. extracted the EMT of the pion from the process $\gamma^{*}\gamma\to\pi^{0}\pi^{0}$ measured on Belle [5]. Burkert et al. obtained the stress inside the proton using deeply virtual Compton scattering data and deeply vector meson production data collected from Jefferson Lab [6, 7, 8, 9, 10]. Duran et. al. accessed the gluon contributions for the proton using near-threshold electroproduction of $J/\psi$ [11]. The precision of these experiments is expected to dramatically improve in the forthcoming era of electron-ion colliders [12, 13, 14]. Lattice simulation of the nucleon and the pion EMT are recently available with a pion mass $M_{\pi}=170\,\mathrm{MeV}$ [15, 16] (cf. [17, 18, 19, 20]). The hadronic EMT were also investigated in various theories and phenomenological models, e.g., [21, 22, 23, 24, 25, 26, 27, 28, 29, 172, 31, 32, 33, 34, 35, 36, 37, 38, 176, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 45, 179, 54, 55]. See Refs. [56, 57] for recent reviews.
@@
One of the challenges to access the hadronic EMT is that $T^{\mu\nu}$ contains the interaction, which is strong at low energy. In this work, we present the first calculation of the charmonium EMT. Charmonium, first discovered 50 years ago in the November revolution, is an ideal system to probe the properties of the strong force. This system consists of a pair of charm and anti-charm quarks. The relatively large charm quark mass $m_{c}\gg\Lambda_{\text{QCD}}$ makes the perturbative QCD description of the quark-antiquark interaction possible. On the other hand, with the binding energy $\alpha_{s}m_{c}\gtrsim\Lambda_{\text{QCD}}$, a non-perturbative treatment is required to describe the properties of the QCD bound states including the stress within. Hence, this bound system involves the interplay of the perturbative and non-perturbative physics of QCD. It is sometimes dubbed as the “hydrogen atom of QCD” [58].
@@
Our method is based on basis light-front quantization (BLFQ) [59]. This method stems from the light-front Hamiltonian formalism, which finds its roots from two lineages [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. One is Dirac [70] and Weinberg’s [71] efforts to simplify relativistic dynamics, which leads to the discovery of the front form of quantum field theory [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. The other is the light-cone current algebra [86, 87, 88, 89, 90, 91] and Feynman’s parton model [92, 93, 94, 95, 96] emerging in deep inelastic scattering [97, 98]. These two complementary perspectives are unified in the light-front Hamiltonian formalism, which allows us to solve QCD as a quantum many-body problem, as described by discretized light-cone quantization (DLCQ) [99] and BLFQ [59].
@@
BLFQ was successfully applied to a number of systems including charmonium, and to a number of hadronic observables including mass spectra, radiative transitions, parton distributions and transverse momentum distributions [101, 102, 103, 104, 105, 106, 107, 108, 109, 52, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 129, 135, 136, 137, 138, 139, 140, 141, 142]. For charmonium, the effective Hamiltonian combines confinement from holographic light-front QCD [143] for the long-distance physics as well as one-gluon exchange interaction from light-front QCD [103] for the short-distance physics [110]. The computed mass spectrum is shown to be within 40 MeV deviation with the PDG compilation [111, 115, 117, 118]. The obtained light-front wave functions (LFWFs) were used to investigate a variety of observables, including the decay constants [110, 111], electromagnetic form factors [114], parton distribution functions [116], and GPDs [114]. In particular, the parameter-free predictions of the radiative transition widths and radiative transition form factors are in remarkable agreement with the experimental measurements whenever available [119, 120, 121].
@@ III Energy-momentum tensor
The 2D Fourier transform of the HMEs in the Drell-Yan frame $q^{+}=0$ can be interpreted as the hadronic EMT on the transverse plane [157, 158, 159, 160, 161, 49, 162],
@@
Note that $t^{\alpha\beta}$ hence $\mathcal{T}^{\alpha\beta}$ depends on the c.m. momentum $P^{\mu}$ of the hadron [159]. The density of the light-front longitudinal momentum $P^{+}$ is,
@@
can be interpreted as the matter density. In Ref. [161], this density is called the internal light-front longitudinal momentum density. We emphasize that $\mathcal{A}(r_{\perp})$ describes the convective distribution of all three momenta. Note that for spin-$\frac{1}{2}$ particles, such as the nucleon, there is an extra contribution from spin,
$\mathcal{T}^{+i}_{ss}(\vec{r}_{\perp};P)=P^{i}_{\perp}\mathcal{A}(r_{\perp})+\big{(}\nabla\times\vec{\mathcal{S}}(r_{\perp})\big{)}^{i}_{\perp}\,.$ (23)
Here, $\vec{\mathcal{S}}(r_{\perp})=2\mathcal{J}(r_{\perp})\vec{s}$ is the spin density, $\vec{s}=s\hat{z}$, and $\mathcal{J}(r_{\perp})$ is the 2D Fourier transform of the GFF $J(Q^{2})$, which exists for spin-$\frac{1}{2}$ particles. N.B. $\mathcal{J}(r_{\perp})$ is normalized to 1/2. The spin current ×$\nabla\times\vec{\mathcal{S}}$ is non-convective.
@@
The r.m.s radius of the matter distribution is,
@@
The first term ${\mathcal{A}(r_{\perp})P^{2}_{\perp}}/{P^{+}}$ represents the light-front energy due to the motion of the particle. The second term in (26) can be interpreted as the distribution of the invariant mass squared $M^{2}$ [49]. Hence, we have abused the notation and introduced the mass squared form factor $M^{2}(q^{2}_{\perp})$ and its Fourier transform, the mass squared density $\mathcal{M}^{2}(r_{\perp})$. In Ref. [161], this density [Eq. (27)] is termed the internal light-front energy density. For spin-$\frac{1}{2}$ particles, the spin term also contributes to the light-front energy density $\mathcal{T}^{+-}(r_{\perp})$,
$\mathcal{T}^{+-}_{ss}(r_{\perp};P)=\\ \frac{P^{2}_{\perp}\mathcal{A}(r_{\perp})+\mathcal{M}^{2}(r_{\perp})+\vec{P}_{\perp}\cdot(\nabla\times\vec{\mathcal{S}}(r_{\perp}))}{P^{+}}\,.$ (29)
@@
This condition is known as the von Laue condition [56]. It is useful to define the r.m.s. radius of the invariant mass squared distribution as,
@@
The above definitions show the advantage of light-front formalism in extracting frame-independent hadronic densities. In general, the hadronic EMT can be parametrized by (the proper) energy density $\mathcal{E}(x)$, pressure $\mathcal{P}(x)$ and (traceless) shear $\Pi^{\alpha\beta}(x)$ as,
@@
This is in contrast to $M^{2}(q^{2}_{\perp})$, the invariant mass squared form factor, in which, the spin contribution is absent. The energy-momentum conservation (24) & (30) [or (31)] implies
@@
i.e. energy density is normalized to the total energy in local rest frame, i.e. the mass. The r.m.s. radius of the energy distribution is,
@@
where, $\hat{r}_{\perp}^{i}=r^{i}_{\perp}/r_{\perp}$. In fact, pressure and shear are the trace and traceless parts of the Cauchy stress tensor $\mathcal{C}^{\alpha\beta}=\Pi^{\alpha\beta}-\mathcal{P}\Delta^{\alpha\beta}$.
Note that the energy density $\mathcal{E}(r_{\perp})$ is related to from the invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ through,
@@
$D\leq\frac{2}{3}M^{2}r^{2}_{A}.$ (46)
Similarly, a related condition, known as the strong energy condition $\mathcal{E}\geq 3\mathcal{P}$ implies, $D\leq({2}/{5})M^{2}r^{2}_{A}$.
@@
$D\equiv D(0)=M\int{\mathrm{d}}^{2}r_{\perp}\,r^{2}_{\perp}\mathcal{P}(r_{\perp}).$ (48)
Polyakov et al. conjectured that $D<0$ for a mechanically stable system [56]. Intuitively, a negative $D$ represents a repulsive core along with an attractive periphery. A negative $D$-term suggests a chain of inequalities,
@@
Figure 1: A 3D picture of the normalized matter density $\mathcal{A}(r_{\perp})$, energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$, scalar density $\theta(r_{\perp})$ and pressure $\mathcal{P}(r_{\perp})$ on the light front. All of these densities, except the pressure, are normalized to unity upon integration over the transverse plane. The pressure $\mathcal{P}$ is normalized to the hadron mass $M$.
The layered picture is consistent with the physics of QCD. From the LFWF representation in Sect. IV, partons with large-$x$ contribute most to the matter density $\mathcal{A}(r_{\perp})$. Therefore, at the core of a charmonium are the valence quarks. For the mass squared density $\mathcal{M}^{2}(r_{\perp})$, small-$x$ partons, known as “wee partons”, also have a significant contribution since the light-front energy $p^{-}\propto(p^{2}_{\perp}+m^{2})/xP^{+}$. The fact that $r_{M^{2}}>r_{A}$ implies that the distribution of the wee partons is broader than the valence quarks. The scalar trace $T^{\alpha}_{\;\;\alpha}$ contains the anomalous gluon fields, whose radius is largest. Therefore, at the outermost of a charmonium, there are glueballs and meson clouds. Note that the quark and the antiquark in our model are effective degrees of freedom. Even though there is no explicit gluons, some effects of them have been incorporated.
In the literature, it is popular to consider the decomposition of the hadron mass into contributions from different species (quarks & gluon), with each contribution defined from a gauge-invariant operator. Since mass in relativity is not additive, what actually gets decomposed is the proper energy $E$. To remove the kinetic energy contributions, each species is assumed to be stationary. Physically, this decomposition is in analogy to the hydrostatics of coupled multifluids [164, 165]. In principle, the invariant mass squared can be decomposed in a similar fashion, except that there is no need to assume that the hadron or its components are at rest since the invariant mass squared $M^{2}$ is a Lorentz scalar. In the previous work [49], some of us proposed to decompose the invariant mass squared $M^{2}$ into a free (or kinetic) part and an interaction (or potential energy) part. This decomposition is more familiar in classical and quantum mechanics although it is not gauge invariant as is the case of the Coulomb potential. Nevertheless, it provides insights into the strong force within charmonium, as we will show in Sect. V.

### v2

@@ Abstract
We investigate the gravitational form factors of charmonium. Our method is based on a Hamiltonian formalism on the light front known as basis light-front quantization. The charmonium mass spectrum and light-front wave functions were obtained from diagonalizing an effective Hamiltonian that incorporates confinement from holographic QCD and one-gluon exchange interaction from light-front QCD. We proposed a quantum many-body approach to construct the hadronic matrix elements of the energy momentum tensor $T^{++}$ and $T^{+-}$, which are used to extract the gravitational form factors $A(Q^{2})$ and $D(Q^{2})$. The obtained form factors satisfy the known constraints, e.g. von Laue condition. From these quantities, we also extract the energy, pressure and light-front energy distributions of the system. We find that hadrons are multi-layer systems.
@@ I Introduction
The year 2023 marked the 50 year anniversary of quantum chromodynamics (QCD). However, our understanding of the strong interaction remains incomplete [1]. One of the challenges is the strong-coupling nature of QCD. As such, the strong force within hadrons is dramatically different from what the QCD Lagrangian naïvely tells us, in contrast to weakly coupled systems, e.g. the hydrogen atom. The primary operators that describe the distribution of stress within a system are the energy-momentum tensor (EMT) $T^{\mu\nu}$ [2]. These operators also dictate how the system gravitates, and conversely, how it responds to the gravitational field [3]. The access to the hadronic matrix elements (HME) of the EMT has been a hot topic in recent literature. Kumano et al. extracted the EMT of the pion from the process $\gamma^{*}\gamma\to\pi^{0}\pi^{0}$ measured on Belle [5]. Burkert et al. obtained the stress inside the proton using deeply virtual Compton scattering data and deeply vector meson production data collected from Jefferson Lab [6, 7, 8, 9, 10]. Duran et. al. accessed the gluon contributions for the proton using near-threshold electroproduction of $J/\psi$ [11]. The precision of these experiments is expected to dramatically improve in the forthcoming era of electron-ion colliders [12, 13, 14]. Lattice simulation of the nucleon and the pion EMT are recently available with a pion mass $M_{\pi}=170\,\mathrm{MeV}$ [15, 16] (cf. [17, 18, 19, 20]). The hadronic EMT were also investigated in various theories and phenomenological models, e.g., [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 45, 53, 54, 55]. See Refs. [56, 57] for recent reviews.
@@
One of the challenges to access the hadronic EMT is that $T^{\mu\nu}$ contains the interaction, which is strong at low energy. In this work, we present the first calculation of the charmonium EMT. Charmonium, first discovered 50 years ago in the November revolution, is an ideal system to probe the properties of the strong force. This system consists of a pair of charm and anti-charm quarks. Since $m_{c}\gg\Lambda_{\text{QCD}}$ and the strong coupling $\alpha_{s}(m_{c})\ll 1$, phenomena occurring at the scale $\sim m_{c}$ may be treated perturbatively. On the other hand, with the binding energy $\alpha_{s}m_{c}\gtrsim\Lambda_{\text{QCD}}$, a non-perturbative treatment is required to describe the properties of the QCD bound states including the stress within. Hence, this bound system involves the interplay of the perturbative and non-perturbative physics of QCD. It is sometimes dubbed as the “hydrogen atom of QCD” [58].
@@
Our method is based on basis light-front quantization (BLFQ) [59]. This method stems from the light-front Hamiltonian formalism, which finds its roots from two lineages [60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. One is Dirac [70] and Weinberg’s [71] efforts to simplify relativistic dynamics, which leads to the discovery of the front form of quantum field theory [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. The other is the light-cone current algebra [86, 87, 88, 89, 90, 91] and Feynman’s parton model [92, 93, 94, 95, 96] emerging in deep inelastic scattering [97, 98]. These two complementary perspectives are unified in the light-front Hamiltonian formalism, which allows us to tackle QCD as a quantum many-body problem, as described by discretized light-cone quantization (DLCQ) [99] and BLFQ [59].
@@
BLFQ was successfully applied to a number of systems including charmonium, and to a number of hadronic observables including mass spectra, radiative transitions, parton distributions and transverse momentum distributions [100, 101, 102, 103, 104, 105, 106, 107, 108, 52, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 128, 134, 135, 136, 137, 138, 139, 140, 141]. For charmonium, the effective Hamiltonian combines confinement from holographic light-front QCD [142] for the long-distance physics as well as one-gluon exchange interaction from light-front QCD [102] for the short-distance physics [109]. The computed mass spectrum is shown to be within 40 MeV deviation with the PDG compilation [110, 114, 116, 117]. For the ground state $\eta_{c}$, the mass deviation is about $\sim 15\%$ of the total binding energy. The obtained light-front wave functions (LFWFs) were used to investigate a variety of observables, including the decay constants [109, 110], electromagnetic form factors [113], parton distribution functions [115], and GPDs [113]. In particular, the parameter-free predictions of the radiative transition widths and radiative transition form factors are in remarkable agreement with the experimental measurements whenever available [118, 119, 120].
@@ III Energy-momentum tensor
The 2D Fourier transform of the HMEs in the Drell-Yan frame $q^{+}=0$ can be interpreted as the hadronic EMT on the transverse plane [161, 157, 158, 159, 160, 162, 49, 163],
@@
Note that $t^{\alpha\beta}$ (hence $\mathcal{T}^{\alpha\beta}$) depends on the c.m. momentum $P^{\mu}$ of the hadron [159]. The density of the light-front longitudinal momentum $P^{+}$ is [164, 161],
@@
can be interpreted as the matter density. In Refs. [164, 161, 162], this density is called the internal light-front longitudinal momentum density. We emphasize that $\mathcal{A}(r_{\perp})$ describes the convective distribution of all three momenta. Note that for spin-$\frac{1}{2}$ particles, such as the nucleon, there is an extra contribution from spin,
$\mathcal{T}^{+i}_{ss}(\vec{r}_{\perp};P)=P^{i}_{\perp}\mathcal{A}(r_{\perp})+\big{(}\vec{\nabla}\times\vec{\mathcal{S}}(r_{\perp})\big{)}^{i}_{\perp}\,.$ (23)
Here, $\vec{\mathcal{S}}(r_{\perp})=2\mathcal{J}(r_{\perp})\vec{s}$ is the spin density, $\vec{s}=s\hat{z}$, and $\mathcal{J}(r_{\perp})$ is the 2D Fourier transform of the GFF $J(Q^{2})$, which exists for spin-$\frac{1}{2}$ particles. N.B. $\mathcal{J}(r_{\perp})$ is normalized to 1/2. The spin current ×$\vec{\nabla}\times\vec{\mathcal{S}}$ is non-convective.
@@
The 3D r.m.s radius of the matter distribution is,
@@
The first term ${\mathcal{A}(r_{\perp})P^{2}_{\perp}}/{P^{+}}$ represents the light-front energy due to the motion of the particle. Here, we have introduced a new form factor $M^{2}(q^{2}_{\perp})$ and its 2D light-front density $\mathcal{M}^{2}(r_{\perp})$. In Refs. [161, 162], this density [Eq. (27)] is termed the internal light-front energy density. We find it is better to interpret $\mathcal{M}^{2}(r_{\perp})$ as the distribution of the invariant mass squared. Hereafter, we will use these two terms interchangeably. For spin-$\frac{1}{2}$ particles, the spin term also contributes to the light-front energy density $\mathcal{T}^{+-}(r_{\perp})$,
$\mathcal{T}^{+-}_{ss}(r_{\perp};P)=\\ \frac{P^{2}_{\perp}\mathcal{A}(r_{\perp})+\mathcal{M}^{2}(r_{\perp})+\vec{P}_{\perp}\cdot(\vec{\nabla}\times\vec{\mathcal{S}}(r_{\perp}))}{P^{+}}\,.$ (29)
@@
As we will see later, this forward limit constraint, also known as the von Laue condition, is a necessary condition for the force equilibrium of a hadron [27, 56]. It is useful to define the r.m.s. radius of the invariant mass squared distribution as,
@@
The above definitions show the advantage of light-front formalism in extracting frame-independent hadronic densities. In general, the hadronic EMT can be parametrized by (the proper) energy density $\mathcal{E}(x)$, pressure $\mathcal{P}(x)$ and (traceless) shear $\Pi^{\alpha\beta}(x)$ as [49, 166],
@@
This is in contrast to $M^{2}(q^{2}_{\perp})$, the light-front energy form factor, in which, the spin contribution is absent. The energy-momentum conservation (24) & (30) [or (31)] implies
@@
i.e. energy density is normalized to the total energy in local rest frame, i.e. the mass. The 3D r.m.s. radius of the energy distribution is,
@@
where, $\hat{r}_{\perp}^{i}=r^{i}_{\perp}/r_{\perp}$. The shear is “traceless”, $\Delta_{\alpha\beta}\Pi^{\alpha\beta}=0$, where $\Delta^{\alpha\beta}=u^{\alpha}u^{\beta}-g^{\alpha\beta}$ is the spatial metric tensor. In fact, pressure and shear are the trace and traceless parts of the Cauchy stress tensor $\mathcal{C}^{\alpha\beta}=\Pi^{\alpha\beta}-\mathcal{P}\Delta^{\alpha\beta}$.
Note that the energy density $\mathcal{E}(r_{\perp})$ is related to the invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ through,
@@
$D\leq\frac{2}{3}M^{2}r^{2}_{A}-1.$ (46)
These conditions are trivially satisfied for negative $D$, which is conjectured for stable systems such as the proton [167, 168].
@@
$D\equiv D(0)=\frac{3}{2}M\int{\mathrm{d}}^{2}r_{\perp}\,r^{2}_{\perp}\mathcal{P}(r_{\perp}).$ (48)
It is conjectured that $D<0$ for a mechanically stable system [167, 168]. Intuitively, a negative $D$ represents a repulsive core along with an attractive periphery. A negative $D$-term suggests a chain of inequalities,
@@
Figure 1: A 3D picture of the normalized matter density $\mathcal{A}(r_{\perp})$, energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$, aka. internal light-front energy density, scalar density $\theta(r_{\perp})$ and pressure $\mathcal{P}(r_{\perp})$ on the light front. All of these densities, except the pressure, are normalized to unity upon integration over the transverse plane. For the pressure, we plot $\mathcal{P}(r_{\perp})/M$.
The layered picture is consistent with the physics of QCD. From the LFWF representation in Sect. IV, partons with large-$x$ contribute most to the matter density $\mathcal{A}(r_{\perp})$. Therefore, at the core of a charmonium are the valence quarks. For the invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$, small-$x$ partons, known as “wee partons”, also have a significant contribution since the light-front energy $p^{-}\propto(p^{2}_{\perp}+m^{2})/xP^{+}$. The fact that $r_{M^{2}}>r_{A}$ implies that the distribution of the wee partons is broader than the valence quarks. The scalar trace $T^{\alpha}_{\;\;\alpha}$ contains the anomalous gluon fields, whose radius is largest. Therefore, at the outermost of a charmonium, there are glueballs and meson clouds. Note that the quark and the antiquark in our model are effective degrees of freedom. Even though there is no explicit gluons, some effects of them have been incorporated.
In the literature, it is popular to consider the decomposition of the hadron mass into contributions from different species (quarks & gluon), with each contribution defined from a gauge-invariant operator. Since mass in relativity is not additive, what actually gets decomposed is the proper energy $E$. To remove the kinetic energy contributions, each species is assumed to be stationary. Physically, this decomposition is in analogy to the hydrostatics of coupled multifluids [169, 170]. In principle, the light-front energy can be decomposed in a similar fashion, except that there is no need to assume that the hadron or its components are at rest since the total internal light-front energy $M^{2}$ is a Lorentz scalar. In the previous work [49], some of us proposed to decompose the light-front energy into a free (or kinetic) part and an interaction (or potential energy) part. This decomposition is more familiar in classical and quantum mechanics although it is not gauge invariant as is the case of the Coulomb potential. Nevertheless, it provides insights into the strong force within charmonium, as we will show in Sect. V.